All ISS systems continue to function nominally, except those noted previously or below. Saturday - off duty day for CDR Volkov, FE-1 Kononenko & FE-2 Chamitoff.
ATV Adieu: At 5:31pm EDT last evcening, ATV1 (Automated Transfer Vehicle 1) "Jules Verne" successfully undocked from the ISS Service Module (SM) aft port, with Sergei Volkov and Oleg Kononenko standing by to monitor separation maneuvers and telemetry parameters. As planned, Kononenko recorded imagery of the ATV front cone during departure. As ATV2 is being assembled at Bremen/Germany, ATV1 performed nominal separation burns and is scheduled for re-entry on 9/29 at night time, to be observed from two high-flying NASA planes.
Due to the late-afternoon ATV undocking, the crew's sleeptime last night began 1.5 hours later than usual (7:00pm). Wake-up this morning was slipped to 3:30am. Tonight the cycle moves back to normal (5:30pm-2:00am EDT).
At ~9:45am, the crewmembers conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-Houston and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.
CDR Volkov downloaded the structural dynamics data collected by the IWIS NCU/RSUs (Internal Wireless Instrumentation System/Network Control Unit/Remote Sensor Units) last evening during the ATV undocking, for subsequent downlink via OCA. Later, Sergey deactivated the IWIS.
In the JAXA Kibo laboratory, FE-2 Chamitoff likewise powered off the MMA (Microgravity Measurement Apparatus) NCU/RSU at the Ryutai rack UDC (Utility DC-to-DC Converter) unit.
After transitioning the JSL (Joint Station LAN) network to the new Netgear wireless APs (Access Points, WAPs) which provide the ISS with WiFi (wireless+Ethernet) connectivity, Gregory today repeated functionality tests, abandoned earlier this week, in three Kibo JPM (JEM Pressurized Module) locations from the wireless SSC -11 laptop, and later also in the COL (Columbus Orbital Laboratory). Afterwards switching to "Proxim" APs, the new WiFi "Dolphin" BCRs (Barcode Readers) were also tested.
In the following 3-hr. VolSci (Voluntary Weekend Science) session, Chamitoff focused again on SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellites), performing algorithm/software development for the experiment as a stand-alone activity, i.e., with no ground support required from PD (Payload Developer) or POIC (Payload Operations & Integration Center/Huntsville) personnel.
Kononenko conducted the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]
Working off their discretionary "as time permits" task list, Sergey and Oleg ran another session with the Russian biomedical MBI-15 "Pilot-M"/NEURO signal response experiment, with Volkov today's subject and Kononenko assisting. Afterwards, the Pilot-M & Neurolab-2000M gear was left connected for the upcoming run for Oleg. [MBI-15 requires a table, ankle restraint system, eyeball electrodes for an EOG (electrooculogram), and two hand controllers (RUO & RUD) for testing piloting skill in "flying" simulations on a laptop (RSK1) under stopwatch control, as well as for studying special features of the psychophysiologic response of cosmonauts to the effects of stress factors in flight.]
The station residents conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exercise device (CDR, FE-1, FE-2) and VELO bike with bungee cord load trainer (CDR, FE-1).
Later, Gregory transferred the exercise data files to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).
SSK Update: After collecting the scheduled microbiological air samples with the MAS (Microbial Air Sampler) kit yesterday, microbiological surface sampling with the SSK (Surface Sampling Kit), also scheduled, had to be deferred due to crew time constraints.
S1 Radiator Cover Sheet Debonding Update: Imagery confirms that a face sheet has peeled back on one panel (#7) of the S1 TRRJ (Thermal Radiator Rotary Joint) Radiator 3, for currently unknown reasons. Teams are looking for any changes and/or movement of the face sheet since the damage was discovered on 9/4. Forward plans to assess the root cause are being worked. Current heat rejection requirements are far below the ETCS (External Thermal Control System) heat rejection capability; therefore no operational impacts have been identified from a heat rejection standpoint. [Each of the eight radiator panels has two thermal face sheets, one on each side, which are epoxy-glued and autoclaved to the honeycomb structure of the radiator. The face sheets have a thermal surface coating designed to maximize heat reflection & absorption transfer to space. Further investigation is underway, and specialists are currently keeping a camera on the panel for monitoring.]
Conjunction Update: When a conjunction prediction last night (8:00pm EDT) for a piece of Kosmos 2421 debris showed a miss distance of 0.67 km at a TCA (Time of Closest Approach) of ~10:00am this morning, leaving insufficient time for calculating & executing a DAM (Debris Avoidance Maneuver) with SM propulsion, Safe Haven instructions were uplinked to the crew. Subsequent tracking data reduced the PC (Probability of Collision) drastically, and when it approached zero this morning, "Safe Haven" was called off. [For the Safe Haven scenario, the crew would have relocated the Airlock PCS (Portable Computer System) laptop to the SM, then closed out and egressed the two Kibo modules (JLP, JPM), COL, Node-2, US Lab, Node-1/PMA-1, closing all hatches to the RS (Russian Segment), which contains ECLSS, and then retreating to the Soyuz crew return vehicle to await TCA.]
Weekly Science Update (Expedition Seventeen -- Week 20)
3-D SPACE: In progress.
ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module.
ANITA (Analyzing Interferometer for Ambient Air): Continuing.
BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.
CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.
CFE (Capillary Flow Experiment): Reserve.
CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Samples returned on 1J.
CSI-2/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): In progress.
CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.
CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.
EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.
ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.
EPO (Educational Payload Operations): Reserve.
ETD (Eye Tracking Device): Completed.
EuTEF (European Technology Exposure Facility): Due to safety concerns identified for the PLEGPAY instrument (when operated in Plasma Discharge mode), the entire EuTEF platform was put in survival mode on 9/1 at around 11:00am EDT (just prior to 29P undocking). Since then, the EuTEF power feeder#1 has been de-activated and no science acquisition is possible. Teams on the ground are closely monitoring the situation and assessing science impacts. EuTEF will be unable to stay in this situation for a long period of time. -- DEBIE-2: Inactive; -- DOSTEL: Inactive; -- EuTEMP: Inactive; -- EVC: Inactive; -- EXPOSE: Inactive; -- FIPEX: Inactive; -- MEDET: Inactive; -- PLEGPAY: Inactive; -- TRIBOLAB: Inactive.
FSL (Fluid Science Laboratory): FSL is nominal.
GEOFLOW: On 8/27, the Science Run#4 was interrupted in order to prepare the FSL Facility for the DAM (Debris Avoidance Maneuver) which took place on that day. GEOFLOW Science Run#1 to Run#4 data have been successfully downlinked and handed over to the science team in Germany. We have gotten very promising results for GEOFLOW so far! The FSL Facility Core Element (FCE) was locked on 8/27 and remains locked for the ISS vehicle traffic (29P undock, ATV undock, 30P docking). The FSL FCE will remain locked at least until after the Progress 30P docking planned on 9/12.
HDTV System Test DL (JAXA): The G1 camcorder function check using MPC was partially conducted on 9/05. Since the Earth view from JPM windows was not able to be performed, it will be rescheduled.
IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.
InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.
Integrated Immune: In progress.
KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.
LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Planned.
Marangoni Experiment for ISS (JAXA Fluid Physics Experiment Facility): The second Marangoni experiment cycle is in progress and will be completed on 9/07.
Micro-G Clay (JAXA EPO): The Clay figures were retrieved from the JPM wall on 8/28 and stowed for return on ULF2.
MISSE (Materials ISS Experiment): Ongoing.
Moon Photography from ISS (JAXA EPO): Complete.
MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 "Pirs".
MSG-SAME (Microgravity Science Glovebox): Complete.
NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.
NUTRITION w/REPOSITORY: In progress.
PADLES (Passive Dosimeter for Lifescience Experiment in Space): In progress.
PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): The PCRF reconfiguration was completed nominally on 8/28.
PMDIS (Perceptual Motor Deficits in Space): Complete.
SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.
SHERE (Shear History Extensional Rheology Experiment): "Greg, you have almost completed all of the planned test points for SHERE. We thank you very much for all the fantastic science you have given us and all your hard work. You have given us science far beyond the original plan, and your results have sparked much discussion and excitement among the team. Thank you! We will send you more results as we analyze the data, and we hope to work with you again soon."
SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): "Greg, thanks for completing your Actiwatch download activity. Your data continues to look great. To avoid the no communication error you got, the stowage notes for the next Sleep activity will direct you to another Actiwatch Reader whose battery power functionality has been verified."
SOLAR (Solar Monitoring Observatory): The current Sun visibility window opened on 8/25, but unfortunately due to high values of ISS Yaw, Pitch and Roll, it was not possible to track the Sun before 8/28. The science acquisition is currently taking place without any problems. SOLAR performed two crisscross measurements on 8/29 and on 9/01. The present Sun observation window should normally close on 9/06.-- SOVIM: continuously acquiring science; -- SOLSPEC: acquiring science, daily calibration and Sun spectrum measurements; -- SOLACES: acquiring science, regular command scripts performed.
SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): "Greg, thanks for supporting SPHERES for Voluntary Science. Test #12 went well!"
Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.
TRAC (Test of Reaction & Adaptation Capabilities): Planned.
WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Teams on the ground continue to assess the results of the BIOLAB Rotor A Bellow Test, which gave faulty signals for 4 out of 6 Reference ECs (Experiment Containers). The next troubleshooting step is currently planned on 9/5 with the Rotor B Actuator test and Rotor B bellow test, all performed from ground.
CEO (Crew Earth Observations): Through 9/04 the ground has received a total of 5,796 frames of CEO images for review and cataloging. "Photos acquired with times corresponding to our CEO target list request times are reviewed first and this week included: Santorini Volcanic Complex, Greece (acquired and under review); Kwanza Basin Angola (under review-more clouds and smoke than we expected); Hurricane Gustav (acquired, confirmed, and published by PAO); High Central Andean Glaciers (some acquired-under review); Cairo Egypt (not acquired); Yellowstone National Park (useful imagery acquired-under review); and Hurricane Ike (acquired, confirmed, and published by PAO). A good week's work! Thank you for the prompt downlink of dynamic events like hurricanes. Your dazzling high-oblique view, through breaks in the clouds, of the middle Amazon River basin will be published on NASA/GSFC's Earth Observatory Site this weekend. Sun glint-enhanced water features provide rare visual insight into the structures and processes of this great river system. Good eye!"
CEO photo targets uplinked for today were Hurricane Ike, Atlantic Ocean (Dynamic Event. Hurricane Ike remains a well-formed storm. It has weakened slightly, but is predicted to regain strength as it continues westwards towards the Bahamas. Looking to the right of track for the storm), Red River Basin, TX (the Red River basin along the Texas/Oklahoma border is the focus of ongoing development, providing an opportunity to track land use/land cover change over time. Overlapping nadir frames - taken along track as ISS approached, crossed, and then departed the river channel almost directly to the north of the Dallas/Forth Worth, TX metro area - were requested), and Jarvis Island, equatorial Pacific (weather is predicted to be clear over Jarvis Island, providing an opportunity for high resolution imagery of the island and surrounding coral reefs. Greg had a nadir pass over the island. Images of the island are useful for mapping shoreline geomorphology and reef extent over time).
CEO photography can be studied at this "Gateway" website:
http://eol.jsc.nasa.gov (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).
Week 21 Scheduled Main Activities:
ISS Orbit (as of this morning, 8:57am EDT [= epoch]):
Mean altitude -- 353.2 km
Apogee height -- 357.9 km
Perigee height -- 348.6 km
Period -- 91.60 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0006956
Solar Beta Angle -- 23.0 deg (magnitude increasing)
Orbits per 24-hr. day -- 15.72
Mean altitude loss in the last 24 hours -- 89 m
Revolutions since FGB/Zarya launch (Nov. 98) -- 56133
Significant Events Ahead (all dates Eastern Time, some changes possible.):
09/08/08 -- Progress M-64/29P de-orbit (~4:45pm)
09/10/08 -- Progress M-65/30P launch (~3:49:45pm)
09/12/08 -- Progress M-65/30P docking (SM aft, ~5:08pm DM)
09/29/08 -- ATV de-orbit (nighttime re-entry for observation from 2 NASA planes; 9:12pm)
10/01/08 -- NASA 50 Years (official)
10/10/08 -- STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4) 12:33am [new target date]
10/11/08 -- Progress M-65/30P undocking (from SM aft)
10/12/08 -- Soyuz TMA-13/17S launch (~3:03am EDT; Lonchakov, Fincke, Garriott)
10/14/08 -- Soyuz TMA-13/17S docking (FGB nadir port, ~4:51am)
10/23/08 -- Soyuz TMA-12/16S undocking (DC1 nadir) or 10/24?
11/12/08 -- STS-126/Endeavour/ULF2 launch - MPLM Leonardo, LMC [new target date]
11/14/08 -- STS-126/Endeavour/ULF2 docking
11/20/08 -- ISS 10 Years
11/25/08 -- Progress M-65/30P undocking & deorbit
11/26/08 -- Progress M-66/31P launch
11/30/08 -- Progress M-66/31P docking
02/09/09 -- Progress M-66/31P undocking & deorbit
02/10/09 -- Progress M-67/32P launch
02/12/09 -- Progress M-67/32P docking
02/12/09 -- STS-119/Discovery/15A launch - S6 truss segment
02/14/09 -- STS-119/Discovery/15A docking
02/24/09 -- STS-119/Discovery/15A undocking
02/26/09 -- STS-119/Discovery/15A landing (nominal)
03/25/09 -- Soyuz TMA-14/18S launch
03/27/09 - Soyuz TMA-14/18S docking (DC1)
04/05/09 -- Soyuz TMA-13/17S undocking
04/07/09 -- Progress M-67/32P undocking & deorbit
05/15/09 -- STS-127/Endeavour/2J/A launch - JEM EF, ELM-ES, ICC-VLD
05/25/09 -- Soyuz TMA-15/19S launch
05/27/09 -- Six-person crew on ISS (following Soyuz 19S docking)
07/30/09 -- STS-128/Atlantis/17A - MPLM(P), last crew rotation
10/15/09 -- STS-129/Discovery/ULF3 - ELC1, ELC2
12/10/09 -- STS-130/Endeavour/20A - Node-3 + Cupola
02/11/10 -- STS-131/Atlantis/19A - MPLM(P)
04/08/10 -- STS-132/Discovery/ULF4 - ICC-VLD, MRM1
05/31/10 -- STS-133/Endeavour/ULF5 - ELC3, ELC4 (contingency).